DIFFERENTIAL GEOMETRY/PDE SEMINAR

Wednesday, March 28 Padelford C-36 3:45 PM

Geometrical completely integrable systems and Hamiltonian stationary Lagrangian surfaces

Frédéric Hélein (CMLA, ENS de Cachan, France)

In the past 20 years, many classical problems in the differential geometry of surfaces have been understood as completely integrable systems. We review these developments and present some new examples of the same type.